### v1

# Sequential Quadratic Sum-of-squares Programming for Nonlinear Control Systems

Abstract
Many problems in nonlinear systems analysis and control design, such as local region-of-attraction estimation, inner-approximations of reachable sets or control design under state and control constraints can be formulated as nonconvex sum-of-squares programs. Yet tractable and efficient solution methods are still lacking, limiting their application in control engineering. To address this gap, we propose a filter line-search algorithm that solves a sequence of quadratic subproblems. Numerical benchmarks demonstrate that the algorithm can significantly reduce the number of iterations, resulting in a substantial decrease in computation time compared to established methods for nonconvex sum-of-squares programs. An open-source implementation of the algorithm along with the numerical benchmarks is provided.
Index Terms:
Constraint control, Lyapunov methods, Numerical algorithms, Optimization algorithms, Sum-of-squares optimization, Variational methods
I Introduction
Sum-of-squares (SOS) optimization provides a computationally tractable framework for certifying polynomial nonnegativity [42]. For nonlinear systems with polynomial dynamics, this property is powerful as many problems arising in nonlinear system analysis and control design can be reformulated into the question of nonnegativity by the Positivstellensatz [47] and thus solved by SOS programming. For convex problems, for example, certification of global asymptotic stability of nonlinear systems, the SOS problem can be transcribed into and solved by semi-definite programs (SDPs). Depending on the problem size, this powerful tool comes with potential high-computational effort. Over the past decades substantial improvement in computational performance can be observed for convex SOS problems, not only due to the improvements in computing hardware. This includes efficient conic solvers such as [4, 22, 38, 21, 23], specialized SOS solver [41], making use of relaxations to (scaled) diagonally dominant cones [2], anytime algorithms [17], efficient toolboxes in different programming languages, among others, [40, 29, 14] or, exploiting the problem structure, such as symmetry [31] or sparsity [1].
However, many practical control engineering problems—such as stability, invariance, and controllability—typically hold only over local regions of the state space [8, 9, 24, 46, 52, 53, 26, 11, 33, 30, 34, 3, 10, 48, 45, 32]. Verifying these properties within such local regions leads to nonconvex SOS optimization problems, which prohibits the direct usage of SDPs. Compared to the convex case, less improvement has been made for the nonconvex cases. State-of-the art methods to deal with the non-convexity are iterative schemes [8, 9, 24, 46, 52, 53, 26, 11, 33, 30, 34, 3, 10, 48, 45, 32]. These schemes include bisection for quasiconvex problems (see [46] for details), coordinate descent, sometimes also referred as alternating-direction, with convex subproblems, and hybrid approaches combining coordinate descent with bisections for quasiconvex subproblems. Table I shows which method is used for different control applications. Coordinate descent iteratively optimizes over one variable at a time while keeping others fixed, simplifying nonconvex problems into sequences of convex SOS programs. However, unless the underlying problem is quasiconvex, these techniques lack convergence guarantees, may require many iterations with unpredictable performance and require a feasible initial guess. Additionally, more effort is needed by the user to split the nonconvex problem into the convex subproblems. There also exist nonlinear SDP solver, e.g., [19], yet their development seems to be stalled and hence is not further considered.
TABLE I: Overview for different control applications and their solution method using either coordinate-descent for convex subproblems (C-CD) or a hybrid approaches (h-CD) to deal with quasiconvex subproblems.
Application Ref. H-CD C-CD
Region-of-attraction [8, 9, 24, 32]
Reachability [52, 53, 26]
[11, 33, 30]
Control-synthesis [26, 34]
Control Barrier Function [3, 10]
Local Control Lyapunov Function [48]
Control Barrier Function & [45]
Control Lyapunov Function
Nonlinear Observer [48]
Inspired by sequential convex programming, we proposed sequential (linear) SOS programming [12]. This method is a promising candidate to reduce the number of iterations compared to coordinate descent and provides (local) convergence guarantees. However, it lacks a comprehensive globalization strategy, does not explicitly address constraint violations, and it lacks an available efficient implementation for practical usage. Furthermore, for practical problems faster convergence is desirable. Thus, leveraging a sequential quadratic programming approach with a suitable globalization strategy is desirable.
Efficient constraint violation checks are crucial, particularly in nonlinear optimization. While simple function evaluations in combination with some norm are used in parameter optimization, conic and SOS programming require projections onto the corresponding cone(s) [7]. As no analytical projection operator is available for the SOS cone, projections must be computed using reliable numerical methods
Globalization strategies are critical for ensuring global convergence in Newton’s method [5, Ch.1]. Merit functions, such as $l_{p}$-norm penalty or augmented Lagrangian, balance cost reduction and constraint violation but require careful penalty parameter selection. This can introduce additional computational costs due to the need for iterative penalty adjustments (cf. [36, 6]) and introduce numerical instability if wrongly selected. This issue is especially relevant in the context of SOS constraint violations, as frequent checks are needed for the penalty adjustment, which may significantly increase the computational burden. Trust-region methods offer superior convergence properties [6], yet, they come at a higher computational cost and lead to larger problems due to additional (norm) constraints.
The filter algorithm [20] offers a promising alternative by decoupling feasibility and optimality. It tracks objective values and constraint violations, making it robust to penalty parameter sensitivity and gives more flexibility in selecting the next solution candidate [6]. Unlike Merit functions, the filter explicitly handles infeasibility through a restoration phase [20, Section 3.3], alleviating numerical instability and ensuring global convergence at the cost of a higher implementation effort.
Given these considerations, a filter line-search algorithm appears to be a suitable candidate for the SOS context. We therefore propose such an approach, which solves a sequence of quadratic subproblems and thereby significantly enhances the method introduced in [12]. Our work builds upon an established filter line-search algorithm proposed in [50] and also takes inspiration from [51]. Whereas global convergence for an active set method is proven in [50], this cannot be extended to SOS in a straightforward manner. We leave the global convergence proof for future work and concentrate on the practical aspects and provide a local convergence analysis.
The key contributions of this paper are as follows.
Convergence Analysis
A local convergence analysis for the sequential SOS algorithm with quadratic subproblems is provided.
Design Aspects
SOS specific design aspects for nonlinear programming are investigated and discussed including the efficient constraint violation check and the design of a feasibility restoration phase.
Practical Implementation
An open-source implementation of the proposed filter line-search algorithm is provided.
In several case studies for real-world inspired examples, we demonstrate the effectiveness and practical applicability of the proposed approach for various control problems. Additionally, we compare it to state-of-the-art methods for nonconvex SOS programming.
II Problem Statement
Motivational Example
We start with a practical problem arising in control engineering. Assume a continuous-time nonlinear dynamic system defined by a polynomial vector field $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}[x]^{n}$, where $x\in\mathbb{R}^{n}$ denotes the state vector and $u\in\mathbb{R}^{m}$ is the control input, satisfying $f(0,0)=0$. We are interested to synthesize a (linear) control law $\kappa(x)$ which asymptotically stabilizes the system’s origin in a domain $\mathcal{X}=\{x\in\mathbb{R}^{n}\mid g(x)\leq 0\}$, i.e., we are interested in a non-conservative state-constrained region-of-attraction. We therefore want to maximize the volume of the sublevel set of the ROA while fulfilling two set-containment constraints, i.e.,
$\max\quad$ $\mathrm{vol}(V,\gamma)$
s.t. $\{x\in\mathbb{R}^{n}\mid V\leq\gamma\}\subseteq\{x\in\mathbb{R}^{n}\mid g(x)\leq 0\}$
$\{x\in\mathbb{R}^{n}\mid V\leq\gamma\}\subseteq\{x\in\mathbb{R}^{n}\mid\langle\nabla V,f_{\kappa}\rangle<0\}$
where $\mathrm{vol}(V,\gamma)$ is the Lebesque measure of the $\gamma$-sublevel set of $V$, $f_{\kappa}=f(\cdot,\kappa(\cdot))$ are the closed-loop polynomial dynamics, $\langle\cdot,\cdot\rangle$ denotes the inner product, $V>0$ is the sought Lyapunov function, and $\gamma$ is a stable level of $V$. Statements of forward invariance, asymptotic stability, regions of attraction, passivity, bounds on the $L_{2}$-gain, input-to-state stability, control Lyapunov and barrier functions, and many more properties of nonlinear control systems can be written as dissipation inequalties [18]. Sum-of-squares constraints provide systematic formulations of dissipation inequalities for polynomial dynamical systems.
II-A Sum-of-squares Polynomials
Let $x=(x_{1},\ldots,x_{n})$ be a tuple of indeterminate variables and $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{N}_{0}^{n}$ a multi-index. A polynomial $p$ in $x$ up to degree $d$ is a linear combination
$p=\sum_{|\alpha|_{1}\leq d}c_{\alpha}x^{\alpha}$
where $x^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$ and $\|\alpha\|_{1}=\sum_{i}\alpha_{i}$. Denote $\mathbb{R}[x]$ as the set of polynomials in $x$ with real coefficients $c_{\alpha}\in\mathbb{R}$ up to degree $d$.
Definition 1.
A polynomial $p\in\mathbb{R}_{2d}[x]$ is a sum-of-squares polynomial ($p\in\Sigma_{2d}[x]$) if and only if there exist $m\in\mathbb{N}$ and $p_{1},\ldots,p_{m}\in\mathbb{R}_{d}[x]$ such that $p=\sum_{i=1}^{m}(p_{i})^{2}$.
The cone of sum-of-squares $\Sigma_{d}[x]$ forms a convex cone in $\mathbb{R}_{d}[x]$ with its dual cone $\Sigma_{d}[x]^{*}$, which is isometric to the cone of sum-of-squares polynomials [28].
As stated in [42, Theorem 3.3] a convex sum-of-squares optimization problem can be reduced to an SDP. Given $p\in\mathbb{R}_{d}[x]$, $s_{d}\in\mathbb{N}$ denotes the number of monomials up to degree $d$ of $p$. A polynomial $p\in\mathbb{R}_{2d}[x]$ is sum-of-squares if and only if there exists a matrix $Q\in\mathbb{S}_{s_{d}}^{+}$ satisfying $p=\zeta^{\top}Q\zeta$, where $\zeta\in\mathbb{R}_{d}[x]^{s_{d}}$ is the vector of monomials up to degree $d$, and $\mathbb{S}_{s_{d}}^{+}$ denotes the set of positive semidefinite matrices.
To transcribe the high-level optimization problem into a SOS problem we make use of the so-called generalized $\mathcal{S}$-procedure, a relaxed version of the Positivstellensatz [48].
Theorem 1 ([48]).
Given $p_{0},p_{1},\ldots,p_{N}\in\mathbb{R}[x]$, if there exist $s_{1},\ldots,s_{N}\in\Sigma[x]$ such that $p_{0}-\sum_{k=1}^{N}s_{k}p_{k}\in\Sigma[x]$, then
$\bigcap_{k=1}^{N}\{x\in\mathbb{R}^{n}\mid p_{k}(x)\geq 0\}\subseteq\{x\in\mathbb{R}^{n}\mid p_{0}(x)\geq 0\}$
holds.
II-B Nonconvex Sum-of-squares Optimization
Applying Theorem 1 to the motivational example yields
$\max_{\begin{subarray}{c}V,\kappa\in\mathbb{R}[x]\\ s_{1},s_{2}\in\Sigma[x]\end{subarray}}$ $\mathrm{vol}(V,\gamma)$
s.t. $s_{1}(V-\gamma)-g(x)$ $\in\Sigma[x]$
$s_{2}(V-\gamma)-\langle\nabla V,f_{\kappa}\rangle-\epsilon(x)$ $\in\Sigma[x]$
$V-\epsilon(x)$ $\in\Sigma[x]$
where $s_{1},s_{2}$ are SOS multipliers resulting from Theorem 1. One can observe that the problem is bilinear in the decision variables $s_{1},s_{2},V$ and $\kappa$ is entering nonlinearly into the second constraint. Clearly, this problem cannot be directly transcribed into and solved by an SDP.
We are interested to efficiently solve general nonconvex, possibly nonlinear SOS problems of the form
$\min_{\xi\in\mathbb{R}_{2d}[x]^{n}}f(\xi)\quad\text{s.t. $\xi\in\Sigma_{2d}[x]^{n}$ and $g(\xi)\in\Sigma_{2d^{\prime}}[x]^{m}$}$ (1)
where $f:\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}$ and $g:\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}_{2d^{\prime}}[x]^{m}$ are differentiable functions. Since the degrees $d,d^{\prime}\in\mathbb{N}_{0}$ are fixed, $f$ and $g$ are mappings between finite vector spaces and the gradients $\nabla f(\cdot):\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}$ and $\nabla g(\cdot):\mathbb{R}_{2d}[x]^{n}\to\mathbb{R}_{2d^{\prime}}[x]^{m}$ are finite linear operators. Define the Lagrangian as $L(\xi,\lambda)=f(\xi)-\langle\lambda,g(\xi)\rangle$} with Lagrange variable $\lambda\in(\Sigma_{2d^{\prime}}[x]^{m})^{*}$. Under a suitable constraint qualification, the optimal solution $\xi_{\star}$ of (1) satisfies the Karush–Kuhn–Tucker (KKT) necessary conditions
[…]
where $\mu_{\star}\in(\Sigma_{2d}[x]^{n})^{*}$ is a normal vector. Eq. (2) can compactly be written as the generalized equation
$\varphi(\xi_{\star},\lambda_{\star})+\mathcal{N}(\xi_{\star},\lambda_{\star})\ni 0$
with
$\varphi:(\xi,\lambda)\mapsto(\nabla f(\xi)-\nabla g(\xi)^{*}\lambda,\,g(\xi))$
where $\mathcal{N}(\xi,\lambda)$ denotes the normal cone to $\Sigma_{2d}[x]^{n}\times(\Sigma_{2d^{\prime}}[x]^{m})^{*}$ at $(\xi,\lambda)$.
III Methodology
This section provides the general concept and core idea of the proposed approach. We start with the basic principle behind sequential quadratic (SOS) programming. Unlike [12], we consider quadratic instead of linear subproblems. A local convergence analysis for quadratic problems is provided. Finally, the concept of filter line-search as globalization strategy is introduced which serves as basis for the practical implementation.
III-A Sequential Quadratic SOS Programming
We solve the nonlinear optimization problem (1) by a sequence of local convex sum-of-squares problems relative to a previous candidate solution (or initial guess) $\xi^{k}$. The solution of the local problem is then used to update the solution candidate and another local problem is solved, as shown in Alg. 1. This framework is a standard approach in nonlinear optimization [36].
Input: Initial guess $\xi^{0}$
1 for $k=0,1,\ldots,N^{\mathrm{max}}$ do
2   Solve convex SOS problem, parametrized in $\xi^{k}$, for search direction $\omega^{k}$
3   Perform line search along $\xi^{k}+\alpha\,\omega^{k}$ for step length $\alpha^{k}$
4   Compute correction or restore feasibility, if necessary
5   Set new solution candidate to $\xi^{k+1}:=\xi^{k}+\alpha^{k}\,\omega^{k}$
6 if termination criteria is satisfied then
7    return
8   end if
9  Update Hessian approximation
10 end for
Algorithm 1 Sequential SOS Algorithm
At the iterate $\xi^{k}$, we solve the quadratic problem
$\min_{\xi\in\mathbb{R}_{2d}[x]^{n}}\tfrac{1}{2}$ $\langle H_{k}(\xi-\xi^{k}),\xi-\xi^{k}\rangle+\langle\nabla f(\xi^{k}),\xi-\xi^{k}\rangle$ (3a)
s.t. $\quad g(\xi^{k})+\nabla g(\xi^{k})(\xi-\xi^{k})\in\Sigma_{2d^{\prime}}[x]^{m}$ (3b)
$\quad\xi\in\Sigma_{2d}[x]^{n}$ (3c)
where $H_{k}:\mathbb{R}_{2d}[x]^{n}\to(\mathbb{R}_{2d}[x]^{n})^{*}$ is an approximation of the Hessian of the Lagrangian in (2). Some conic solvers [38, 22] can handle a quadratic cost directly, otherwise one rewrites the quadratic cost into a second-order cone constraint using the epigraph reformulation [7].
The KKT conditions for the quadratic subproblem (3) read
[…]
where $\xi_{+}$, $\lambda_{+}$ and $\mu_{+}$ denote the optimal solution, corresponding Lagrange multipliers and normal vector of (3), respectively. This approach is also known as quasi-Newton method [25] and it exhibits fast local convergence with a well-chosen initial guess; however, it often requires a globalization technique for robust performance from arbitrary starting points. We are going to utilize a line search to compute the new candidate solution as
$\xi^{k+1}$ $=\xi^{k}$ $+\alpha^{k}\underbrace{(\xi_{+}-\xi^{k})}_{\omega^{k}}$ (5a)
$\lambda^{k+1}$ $=\lambda^{k}$ $+\alpha^{k}(\lambda_{+}-\lambda^{k})$ (5b)
where $\alpha^{k}\in(0,1]$ is the chosen step length at iteration $k$. Approximation methods for the Hessian are discussed in Subsection IV-C1.
III-B Local Convergence Analysis
Should we explain, why we only consider the best case, i.e., neglecting any kind of globalization? How is without globalization the ‘best’ case? We should state though, preferably already in the introduction, that the proof by Wächter and Biegler considers active sets methods and cannot be applied to SOS; and that we leave the global convergence proof to future work.
We provide a brief, local convergence analysis of the sequential algorithm without line search ($\alpha^{k}\equiv 1$). If $(\xi_{+},\lambda_{+})$ solves the KKT conditions (4) with $\mu_{+}$, the update (5) can be written as
$(\xi^{k+1},\lambda^{k+1})\in(\xi^{k},\lambda^{k})-(\Phi(\xi^{k},H_{k})+\mathcal{N})^{-1}\big[\varphi(\xi^{k},\lambda^{k})\big]$ (6)
where $(\cdot)^{-1}$ is the (possibly multi-valued) inverse mapping, with
$\Phi(\xi^{k},H_{k}):(\xi,\lambda)$ $\mapsto(H_{k}\xi-\nabla g(\xi^{k})^{*}\lambda,\nabla g(\xi^{k})\xi)$
and candidate solution $\xi^{k}$ and Hessian approximation $H_{k}$. The key observation for the convergence analysis is that (6) constitutes a set-valued Newton update step for the nonlinear KKT conditions (2), where $\Phi(\xi^{k},H_{k})$ is an approximation of the first-order derivative of $\varphi$.
The convergence properties of set-valued Newton methods are a long studied subject of variational analysis citations. Local convergence can usually be established provided that the inverse of $(\Phi(\xi^{k},H_{k})+\mathcal{N})$ in (6) is well defined and the approximation $\Phi(\xi^{k},H_{k})$ is ‘sufficiently good’ – the better the approximation, the faster the rate of convergence. In particular, we make the following assumptions:
Assumption 1.
The exist constants $\zeta,\ell_{k}>0$ such that
  1. (a)
    the inverse $(\Phi(\xi^{k},H_{k})+\mathcal{N})^{-1}$ is locally single-valued around $(\xi_{\star},\lambda_{\star})$ and locally Lipschitz continuous at 0 with constant $\zeta$ uniformly for $\xi^{k}$ around $\xi_{\star}$ and all $H_{k}$;
  2. (b)
    the mapping $\varphi(\xi^{k},\lambda^{k})+\Phi(\xi^{k},H_{k})(\xi^{\prime}-\xi^{k},\lambda^{\prime}-\lambda^{k})$ is locally Lipschitz continuous relative to $(\xi^{k},\lambda^{k})$ at $(\xi_{\star},\lambda_{\star})$ with constant $\ell_{k}$ uniformly for $(\xi^{\prime},\lambda^{\prime})$ around $(\xi_{\star},\lambda_{\star})$ and all $H_{k}$.
Part (a) of Assumption 1 is called strong regularity of the set-valued mapping $(\Phi(\xi^{k},H_{k})+\mathcal{N})$ and corresponds to unique primal-dual solutions to the convex problem (3) that are continuous under first-order perturbations e.g., [15]. Regularity can be enforced by choosing a positive definite approximation $H_{k}$. The Lipschitz constant in part (b) measures how good the linear operator $\Phi(\xi^{k},H_{k})$ describes the change of $\varphi(\xi^{k},\lambda^{k})$ as $(\xi^{k},\lambda^{k})\to(\xi_{\star},\lambda_{\star})$ and can be thought of as generalization of an error in the derivative of $\varphi$. In particular, Assumption 1(b) is satisfied with $\ell_{k}=0$ if $H_{k}$ is the second-order derivative $\nabla_{\xi\xi}^{2}[f(\xi_{\star})-\langle\lambda_{\star},g(\xi_{\star})\rangle]$.
Theorem 2.
Suppose that Assumption 1 holds at a solution $(\xi_{\star},\lambda_{\star})$ to (2); then the sequence generated by (6) is unique and converges
  1. 1.
    linearly, if $\ell_{k}<\zeta^{-1}$;
  2. 2.
    superlinearly, if $\ell_{k}\searrow 0$;
  3. 3.
    quadratically, if $\ell_{k}\equiv 0$;
to $(\xi_{\star},\lambda_{\star})$ provided that $(\xi^{0},\lambda^{0})$ is sufficiently close.
Proof.
By virtue of [16, Theorem 8.5] with $h:(\xi,\lambda)\mapsto\varphi(\xi_{\star},\lambda_{\star})+\Phi(\xi_{\star},H_{k})(\xi-\xi_{\star},\lambda-\lambda_{\star})$, the sequence $\{(\xi^{k},\lambda^{k})\}_{k\in\mathbb{N}}$ generated by (6) is unique and satisfies
$\|(\xi^{k+1},\lambda^{k+1})-(\xi_{\star},\lambda_{\star})\|\leq\zeta\ell_{k}\|(\xi^{k},\lambda^{k})-(\xi_{\star},\lambda_{\star})\|$
for all $k\in\mathbb{N}$, provided that $(\xi^{k},\lambda^{k})$ is close to $(\xi_{\star},\lambda_{\star})$; this implies (super) linear convergence if $\zeta\ell_{k}<1$ (resp., $\zeta\ell_{k}\searrow 0$), and quadratic convergence if $\zeta\ell_{k}\equiv 0$. ∎
III-C Line-search with Filter
This subsection provides the core components of the filter line-search algorithm, which is primarily based on [6, p. 120]. We outline the core parts to make this paper self-contained. The filter line-search algorithm is summarized in Algorithm 2. In general three major steps are conducted: 1) check filter acceptance, 2) check sufficient decrease in cost and 3) augment filter, if necessary.
III-C1 Filter Acceptance
Unlike merit functions, that combine cost $f$ and constraint violation $\theta$ in a single objective, a filter tries to deal with these two conflicting goals in a bi-objective optimization manner. No penalty parameter must be estimated, as is the case for penalty functions or augmented Lagrangian. The filter consists of $l$ so called dominant pairs $\{f(\xi^{l}),\theta(\xi^{l})\in\mathcal{F}$ which form a forbidden region. In a first step, the current trial point is checked for filter acceptance, that is, checking if the trial point lies not in the forbidden region and makes sufficient progress regarding constraint violation. A new trial point, usually the result of (5) for the candidate step length $\alpha$, is accepted to the filter $\mathcal{F}$ if
$f(\xi^{k+1})<f(\xi^{l})\quad\text{or}\\ \quad\theta(\xi^{k+1})<\theta(\xi^{l})\quad\forall(f(\xi^{l}),\theta(\xi^{l}))\in\mathcal{F}$ (7)
Condition (7) alone only provides sufficient reduction of the constraint violation but not for improvement of the cost function. Therefore, additional sufficient decrease conditions for the cost function are checked for an acceptable point, as outlined next. If the candidate fails or violates the first condition (7), a second-order correction may be applied, or the step length is reduced.
III-C2 Sufficient Decrease Conditions
The first condition ensures that the current search direction $d=\xi^{k+1}-\xi^{k}$ is a descent direction
$\nabla f(\xi^{k})^{\top}d<0$ (8)
where $\nabla f(\xi^{k})$ is the gradient of the original cost. A second condition is used and reads
$\alpha(-\nabla f(\xi^{k})^{\top}d)^{s_{\phi}}>\delta\theta(\xi^{k})^{s_{\theta}}$ (9)
where $s_{\phi},s_{\theta}\in(0,1)$ and $\delta\in(0,1)$ are small constants. This condition ensures a sufficiently large progress for the cost function compared to the current constraint violation [6]. Equations (8) and (9) are also referred as f-type switching condition.
In case the f-type switching is fulfilled and additionally if $\theta(\xi^{k+1})<\theta_{\text{min}}$ [51] then the Amijo condition is checked
$\phi(\xi^{k+1})\leq\phi(\xi^{k})+\alpha\rho\nabla\phi(\xi^{k})^{\top}d$ (10)
where $\rho\in[0,\frac{1}{2}]$. Can $\rho$ be zero?
If no reduction in cost is possible (according to f-type) it is checked if either the constraint violation or cost is smaller than the previous iterate with a small envelope
$\theta(\xi^{k+1})\leq\theta(\xi^{k})(1-\gamma_{\theta})\text{ or }\phi(\xi^{k+1})\leq\phi(\xi^{k})-\gamma_{f}\theta(\xi^{k})$ (11)
where $\gamma_{\theta}\in[0,1]$ and $\gamma_{f}\in[0,1]$ are small constants. Can $\gamma$ be zero?
III-C3 Filter Augmentation
Unlike [20], the filter is only augmented if, and only if, any of the sufficient decrease conditions (8), (9) or (10) do not hold. This stricter augmentation rule avoids cycling, that is, the iterates alternately improve one of the two measures, but deteriorate the other [50].
III-C4 Full Algorithm
Algorithm 2 outlines the filter line-search procedure (cf.  [6]). To avoid the Maratos effect [36], we additionally add a second-order correction (details in Section IV-C2). If the second-order correction is able to find an acceptable trial point, the line-search is terminated, otherwise, the step length is reduced. In case the step-length falls below the minimum step-length $\alpha_{\text{min}}$ a feasibility restoration phase is invoked. The primary goal of the feasibility restoration phase is to minimize the constraint violation, such that the trial point is acceptable to the filter.
Input: $\xi^{k},\xi_{+}$
Output: $\xi^{k+1}$
1 Set $\alpha=1$
2 while $\alpha\geq\alpha_{\mathrm{min}}$ do
3   Compute new trial point $\xi_{\alpha}=\alpha(\xi_{+}-\xi^{k})$
4 if Acceptable to filter (7) then
5    if (8) and (9) hold then
6       if (10) holds then
7          return $\xi_{\alpha}$
8
9       else
10            Invoke SOC if $\alpha=1$
11
12         end if
13
14    else
15       if (11) holds then
16          return $\xi_{\alpha}$
17       else
18            Invoke SOC if $\alpha=1$
19
20         end if
21
22      end if
23
24 else
25      Invoke SOC if $\alpha=1$
26
27   end if
28  Adjust step length $\alpha=\frac{1}{2}\alpha$
29
30 end while
31if [(8) and (9)] or (10) do not hold then
32   Augment filter
33 end if
Algorithm 2 Filter Line-search with second-order correction (SOC)
III-D Required SOS Modification
The filter frequently needs to check progress in cost and feasibility. While checking progress in cost is a simple function evaluation, evaluating the constraint violation (feasibility) for SOS constraints is more complicated. In parameter optimization, constraint violation can be assessed by evaluating the constraints at the current iterate and use some norm, e.g., $\theta(\xi^{k}):=\|g(\xi^{k})\|_{\infty}$. For conic optimization, the constraint violation check involves a projection onto the corresponding cone (cf. [7]). Although some cones possess relatively simple (analytical) projection operators, this is not the case for the SOS cone. Since the constraint violation is frequently evaluated, we are interested in precise yet efficient methods for this task. Similarly, a dedicated feasibility restoration,First mention? which primary goal is to reduce constraint violation, must be designed and implemented for SOS.
IV Practical Implementation
[…]
Input: Initial guess $\xi^{0}$, parameters
1 Initialize Hessian $H_{0}$, Initialize filter, $k=0$
2 while $k\leq N_{\mathrm{max}}$ do
3
4 if $\nabla L(\xi^{k},\lambda^{k})||\leq\epsilon_{L}$, and $\theta(\xi^{k})||\leq\epsilon_{\theta}$ then
5      Terminate
6
7 else
8    Solve (3) to obtain $\xi^{*}$ and $\lambda^{*}$
9    if Feasible then
10         Call Algorithm 2 (filter line search)
11       if $\alpha<\alpha_{\mathrm{min}}$ then
12            Call Algorithm 4 (Feas. Restoration) to obtain $\xi_{r}$
13          if $\xi_{r}$ acceptable to filter then
14               Set $\xi^{k}=\xi_{r}$
15               Go to line 3
16          else
17               Terminate (locally infeasible)
18            end if
19
20       else
21            Accept trial point
22         end if
23
24    else
25         Call Algorithm 4 (Feas. Restoration) to obtain $\xi_{r}$
26       if $\xi_{r}$ acceptable to filter then
27            Set $\xi^{k}=\xi_{r}$
28            Go to line 3
29       else
30            Terminate (locally infeasible)
31         end if
32
33      end if
34
35     Compute Hessian approximation
36      set $k=k+1$
37   end if
38
39 end while
Algorithm 3 Sequential Quadratic SOS
IV-A Constraint Violation Estimation
Unlike standard optimization, the constraint violation check for SOS constraints is not a simple function evaluation. In the following, different methods are discussed and compared in a small benchmark.
We assume $n_{g}$ nonlinear constraints. The constraint violation $\theta(\xi^{k})$ for the solution $\xi^{k}$ indicates whether the constraints $g(\xi^{k})\in\Sigma[x]^{n_{g}}$ are satisfied ($\theta(\xi^{k})=0$) or, if not, measures how far ($\theta(\xi^{k})>0$) the solution is from satisfying the constraints. Take $p_{k}:=g(\xi^{k})$. Besides, the last statement is simply not true.
IV-A1 SOS Projection
The projection onto the cone of sum-of-squares can be computed as the convex SOS program
$\theta(\xi^{k}):=\min_{s\in\mathbb{R}[x]^{n_{g}}}\left\|{s}(x)-{p_{k}}(x)\right\|_{2}^{2}\quad\text{s.t. $s\in\Sigma[x]^{n_{g}}$}$ (12)
where the optimal solution for $s$ is the nearest sum-of-squares polynomial to $p_{k}$.
IV-A2 Signed Distance
In [14, Sec. 4.3], a signed-distance metric for generalized cones is introduced. It utilizes a scalarization function [27] to map $p_{k}$ to a finite signed distance between $p$ and the boundary of $\Sigma[x]^{n_{g}}$. The signed-distance metric is defined as
$\mu_{\Sigma,s}:p\mapsto\inf\{r\in\mathbb{R}\mid p+r\odot s\in\Sigma[x]\}$ (13)
where the optimal value for $r$ is the scalar signed distance to the boundary of the cone, $\odot$ denotes the Hadamard product, and $s(x)\in\operatorname*{int}\Sigma[x]$. The SOS optimization problem
$\min_{r\in\mathbb{R}^{n_{g}}}\quad$ $\quad\sum_{j=1}^{n_{g}}r_{j}\quad\text{s.t.}\quad p_{k}(x)+r\odot s(x)\in\Sigma[x]^{n_{g}}$ (14a)
is used to compute the signed-distance metric. We define the overall distance, and hence constraint violation, as
$\theta(x):=\begin{cases}0&\text{if }\mathrm{max}(r^{*})\leq\epsilon\\ \mathrm{max}(r^{*})&\text{otherwise}.\end{cases}$ (15)
where $r^{*}$ is the optimal solution to (14) and $\epsilon>0$ is a tolerance.
IV-A3 Pseudo-Projection via Sampling
Verifying that $p_{k}(x)\in\Sigma[x]^{n_{g}}$ can be replaced by checking the necessary condition $p_{k}(x)\geq 0$ on a finite set of sample points $x^{(i)}\in\mathbb{R}^{n}$, $i=1,\dots,M$, and nonnegativity is checked element-wise. Samples are drawn from a bounded hypercube.
To assess feasibility, the minimum value of each polynomial over the sampled points is computed as
$p_{\min}=\min_{\begin{subarray}{c}j=1,\dots,n_{g}\\ i=1,\dots,M\end{subarray}}p_{k}(x^{(i)})_{j}$ (16)
and the constraint violation is set to
$\theta(\xi^{k}):=\begin{cases}0&\text{if }p_{\min}\geq 0\\ |p_{\min}|&\text{otherwise}.\end{cases}$ (17)
It should be noted that $\theta(\xi^{k})=0$ is neither necessary nor sufficient for $p_{k}\in\Sigma[x]^{n_{g}}$.
Discussion
The SOS projection and signed distance methods both define a distance metric for constraint violation via SDPs. The sampling approach is simple to implement and does not require an SDP solver, however, it is incomplete and its accuracy depends on the number of sample points.
Figure 1 compares the computation times of these methods for increasing numbers $n$ of indeterminate variables, using 10 randomly generated SOS polynomials of degree six. The projection method scales poorly, becoming impractical for larger $n$. In contrast, the signed distance remains efficient, while sampling with 1,000 points is fastest but least accurate.Should we report false-positives for Sampling? Based on this comparison, we implement the signed distance method for its strong balance of accuracy and efficiency.
[…]
Figure 1: Performance comparison of methods used to measure constraint violation in SOS programming.
IV-B Feasibility Restoration
Small step lengths can cause the solver to stall or fail. If $\alpha^{k}<\alpha_{\text{min}}$ or infeasibility is detected in (3), Alg. 3 enters the feasibility restoration phase (lines 10 and 21 in Alg. 3) to maintain robustness [36, p.439]. This phase aims to reduce the constraint violation and provide a solution that is acceptable to the filter of the original problem.
Remark 1.
There is no guarantee that the restoration finds a solution acceptable to the filter. If it does not, we want it to converge to a local (infeasible) point. This information might be used by the user to take corrective action on the problem formulation (cf. [51]).
The feasibility restoration might be invoked after certain progress has already been made towards a local solution, but the constraint violation is still too high. In that case, we want the restoration not to deviate too much from the current, potentially close to (local) optimal solution. To achieve this, we introduce a weighted regularization term based on the current constraint violation [51, 49].
The high-level SOS feasibility restoration problem with regularization reads
$\min_{r\in\mathbb{R}^{n_{g}},\xi_{R}\in\mathbb{R}^{n}}$ $\quad\sum_{i=1}^{n_{g}}r_{i}+\frac{\rho_{k}}{2}\|\xi_{R}-\xi^{k}\|_{2}^{2}$ (18a)
s.t. $\quad g(\xi_{R})+r\odot s\in\Sigma[x]^{n_{g}}$ (18b)
where $\rho_{k}>0$ is a penalty parameter, and $\xi^{k}$ is the iterate at which the restoration phase is initiated. Problem (18) is similar to (14), but it differs in that both the decision variables and the constraint violations are optimized simultaneously. The additional decision variables introduced are negligible compared to the overall problem size. Due to the explicit reduction of constraint violation in the cost function, previous approaches for nonlinear parameter optimization, such as [51], needed a smooth reformulation of the problem using slack variables. In contrast, our approach is already smooth.
The penalty parameter is heuristically determined based on the constraint violation (signed-distance) of the current iterate. In our implementation, we make use of the following power function
$\rho_{k}=\begin{cases}\rho_{\text{max}}&\text{if $\theta(\xi^{k})\leq\epsilon_{\text{feas}}$}\\ \epsilon_{\text{feas}}/\theta(\xi^{k})+\rho_{\text{min}}&\text{else}\end{cases}$ (19)
with $0<\rho_{\text{min}}<\rho_{\text{max}}$ defined by the user. In case of large constraint violation, more weight is put onto feasibility, and vice-versa. The penalty $\rho_{k}$ is computed once at the initiation of the feasibility restoration.
Problem (18) is again nonlinear and hence we apply the previously described filter line-search algorithm to solve this problem. Two modifications are made:
  • The restoration phase does not have its own restoration phase and,
  • the restoration phase is left once an iterate is acceptable to the filter of the original problem with an additional threshold on constraint violation, which takes the form $\theta(\xi_{R}^{k})\leq\eta\theta(\xi^{k})$ with $\eta\in[0,1]$ (cf. [51]).
The adapted algorithm for the feasibility restoration phase is outlined in Alg. 4. In the current implementation, the restoration is not invoked if the constraint violation is already below the assigned threshold.
Input: Initial guess $\xi^{0}$, $s^{0}$, parameters
1 Initialize Hessian $R_{0}$, Augmented filter, $k=0$
2 while $k\leq N_{\mathrm{max}}$ do
3 Solve (3) to obtain $\xi_{R}^{*}$ and $\lambda_{R}^{*}$ and set $\alpha_{k}=1$
4 if Feasible then
5      Execute filter line-search (Algorithm 2)
6    if $\alpha\leq\alpha_{\mathrm{min}}$ then
7         Solver stalled-Terminate
8      end if
9
10 else
11      Problem seems infeasible-Terminate
12   end if
13
14 if Acceptable to augmented filter and $\theta(\xi^{k+1})||\leq\kappa\theta(\xi^{k})$ then
15      Terminate
16
17 else
18      Compute Hessian approximation
19      set $k=k+1$
20   end if
21
22 end while
Algorithm 4 Feasibility Restoration Phase
IV-C Implementation Details
This section provides additional implementation details and practical aspects. This includes Hessian approximation, the details about the implemented second-order correction and parameters used.
IV-C1 Hessian Approximation
In order to obtain an accurate but also positive definite estimate of the Hessian, several methods can be found in the literature. Our implementation provides several options of which the applicability depends, e.g., on the specific problem size. Options include the damped Broyden-Fletcher-Goldfarb-Shanno (BFGS) update [36, Procedure 18.2] and exact Hessian computation in combination with regularization techniques such as the Gershgorin bound [5, Section 2.5], minimum Frobenius norm approach [36, p. 50] or using the mirrored version of the exact Hessian [43]. The initialization of the Hessian is also crucial for convergence of the algorithm. The matrix $H_{0}$ in Alg. 3 is initialized with the identity matrix, unless the user provides a scaling factor to improve convergence (cf. [36, Ch. 6]) or the exact (regularized) Hessian is used.
IV-C2 Second-Order Correction
Merit functions or filters may converge slowly because large steps are often rejected, a phenomenon known as the Maratos effect. A potential cause of slow convergence is an inaccurate linear approximation of the constraint function. To mitigate this, a second-order correction (SOC) step is applied to adjust the search direction and further reduce constraint violations [36]. Quadratic approximations of the constraint functions would address this but lead to a challenging quadratically-constrained quadratic program. An alternative, outlined in [36, p. 544], is employed here. The approach involves replacing the quadratic term in the constraint function’s quadratic approximation with a constant quadratic term based on the current uncorrected search direction $\omega^{k}=(\xi_{+}-\xi^{k})$ (see [36, p. 544] for details). Specifically, the constant term
$g(\xi^{k}+\omega^{k})-\nabla g(\xi^{k})^{\top}\omega^{k}$
is added to (3b) and the adapted quadratic subproblem is then solved to obtain $\xi_{\text{soc}}^{*}$ which is used to adjust the search direction. The thus corrected, new trial point is then given by (with $\alpha=1$)
$\xi^{k+1}=\xi_{\text{soc}}^{*}+\omega^{k}$ (20)
The new trial point from (20) is also checked for filter acceptance. If it is not possible to find a feasible point, the SOC returns unsuccessful and step-length is reduced.
IV-C3 Scaled Termination and Parameter
To improve robustness against numerical errors and poor scaling, we make use of a scaled KKT condition given by [44]
$\left\|\nabla_{\xi}L(\xi^{k},\lambda^{k})\right\|_{\infty}\leq\frac{\epsilon_{\text{opt}}\,\max\left(1,|f(\xi^{k})|\right)+\left\|\langle\lambda^{k},g(\xi^{k})\rangle\right\|_{\infty}}{\left\|\omega^{k}\right\|_{\infty}}$ (21)
that ensures
$|f(\xi^{k+1})-f(\xi^{k})|\leq\epsilon_{\text{opt}}.$ (22)
For a derivation of the scaled KKT condition see [35]. In our implementation, we set the tolerances and parameters of the filter as follows: $\epsilon_{\text{opt}}=10^{-4}$, $\epsilon_{\text{feas}}=10^{-6}$, $s_{f}=2$, $s_{\theta}=0.9$, $\delta=1$ and $\rho=10^{-4}$, $\gamma_{f}=\gamma_{\theta}=10^{-5}$, $\eta=10^{-4}$, $\rho_{\mathrm{min}}=0.01$ and $\rho_{\mathrm{max}}=1$.
V Case Studies
In this section we provide several numerical examples known from nonlinear system analysis and control design. Only nonconvex SOS problems with potentially non-affine system dynamics are considered. The proposed filter line-search algorithm is benchmarked to the state-of-the art method, that is, coordinate-descent (CD) (cf.  [8, 9, 24, 46, 52, 53, 26, 11, 33, 30, 34, 3, 10, 48, 45, 32]). We compare the number of iterations, computation time and cost for given problem size. We denote the problem size by the number of system states $n$, constraints $n_{\text{con.}}$ (both linear and cone constraints) and decision variables $n_{\text{dec.}}$ of the underlying SDP. We report only the problem size of the sequential SOS approach, as it reflects the overall size. In contrast, CD involves multiple smaller subproblems, which may include redundant constraints or variables. A detailed breakdown is available in the supplementary material [39]. This section concludes with a comparison of the main characteristics of the two methods and provides practical aspects such as initial guess generation for the sequential approach.
V-A Preliminaries
[…]
V-B Region-of-Attraction Estimation
We consider the problem to find the region-of-attraction (ROA) for nonlinear closed-loop systems. We are interested to solve the following standard ROA estimation problem
$\min_{\begin{subarray}{c}V\in\mathbb{R}[x],s\in\Sigma[x]\end{subarray}}$ $\quad\|g_{0}-V\|^{2}_{\mathbb{R}[x]}$ (23a)
s.t. $V-\epsilon(x)$ $\in\Sigma[x]$ (23b)
$s(V-\gamma)-\langle\nabla V,f\rangle-\epsilon(x)$ $\in\Sigma[x]$ (23c)
where $V$ is the sought Lyapunov function, $s$ is a SOS multiplier coming from the generalized $\mathcal{S}$-procedure and $f(\cdot)$ is the closed loop dynamic, $g_{0}$ is a pre-defined safe set which we make use of to find a less conservative solution by minimizing the squared $l_{2}$-norm $\|\cdot\|_{\mathbb{R}[x]}$ on the space of polynomials. [^1] A small cost value indicates large portions of the safe set are covered by the ROA function. The stable level set is pre-defined as $\gamma=1$, but could be also a decision variable. Two examples are considered: The first example is taken from [9], where the nonlinear ROA analysis is used to assess the robustness properties of the F/A-18 flight control laws. The second example evaluates scalability and robustness in terms of the number of system states $n$, using an $N$-link (planar) robot arm.
V-B1 F/A-18
We take the seven state closed-loop dynamics from [9, Appendix B]. For both methods, we compute an initial Lyapunov candidate by linearizing the nonlinear closed-loop dynamics. For the sequential approach, we additionally set the initial SOS multiplier $s=x^{\top}x$. Due to infeasibility of CD at the first iteration, we solved the first subproblem via bisection. After the first iteration, we used the pre-defined level set. The results of the two approaches are summarized in Table II.
The sequential approach only needs 19 iterations until convergence, whereas coordinate descent did not converge within the pre-assigned maximum number of iterations. The total time (solver build + solve time) as stated in Table II is substantially less for sequential. The build time for the sequential SOS is $4.81\text{\,}\mathrm{s}$, whereas for coordinate descent, which includes several subproblems, is only $1.15\text{\,}\mathrm{s}$.
TABLE II: Comparison of sequential SOS (S) and coordinate descent (C) for the F-18 ROA problem.
n Size Iter. [-] Total time [s] Cost [-]
$_{\text{con.}} n$$_{\text{dec.}}$ S C S C S C
7 2058 15785 19 100 52.5 277.27 2.42 4557.3
V-B2 N-link Robot Arm
Using an $N$-link (planar) robot arm, the problem scales to $n=2N$ states. For each $N$, the dynamics and a linear quadratic control law (LQR) are computed and a quadratic polynomial approximation for the closed-loop system is calculated. The details are provided in the supplementary material [39]. We solve the problem twice using different initial guesses. In the first case, both algorithms are initialized with the Lyapunov function of the linear system, with the sequential algorithm additionally using a simple quadratic polynomial for the SOS multiplier. In the second, we make use of a clearly infeasible initial guess. A more detailed breakdown is given in Table III and  IV.
Good Initial Guess
The results are summarized in Table III. Both methods achieve optimality (using the assigned thresholds). Final costs are omitted because they are nearly identical, with the sequential method performing slightly better. Notably, the sequential method requires significantly less iterations than coordinate descent—especially for larger number of states—resulting in lower total computation time as shown in Fig. 2 and Table III. In some cases, CD was infeasible in the first iteration in the multiplier subproblem. In these cases we gave CD a grace iteration as indicated by and tried to solve the Lyapunov step. On average, sequential requires 60% less iterations and 54% less time to solve the problems.
Bad Initial Guess
The second case considers the case with a bad initial guess. We initialize all decision variables with negative quadratic polynomials. Such polynomials are clearly infeasible.
Remark 2.
It is unlikely that someone uses such an obviously infeasible initial guess and rather uses an initial guess similar to the first case. This is only made for demonstration purpose. It should be noted that an initial guess like in the first case does not guarantee convergence for coordinate descent in other cases, as demonstrated in the previous F/A-18 example.
Coordinate descent is not able to start the computation because it needs a feasible initial guess. Thus, no results can be reported in this case. In contrast, the proposed sequential algorithm with its feasibility restoration is able to solve the problem to full optimality or in a few cases at least to a feasible solution (indicated by $(\cdot)^{*}$ in Table IV). [^2] Feasible solution means the thresholds for constraint violation are met, but solution might be far from optimal. This clearly shows the enhanced robustness of the proposed approach. The call of the feasibility restoration phase increases the number of iterations significantly for smaller number of states. For medium to large number of states, the number of iterations is only moderately increased. However, the average time per iteration is much larger compared to cases where the feasibility restoration is not called, which is expected.
[…]
Figure 2: Relative performance improvement of sequential SOS (S) with respect to coordinate-descent (CD) in terms of total computation time $T_{(\cdot)}$ and iterations $I_{(\cdot)}$.
TABLE III: Comparison of sequential SOS (S) and coordinate descent (C) for the N-link robotic arm ROA example using a good initial guess.
n Size Iter. [-] Time [s]
$_{\text{con.}} n$$_{\text{dec.}}$ S C S C
44 $85$ $248$ 77 $23$ $0.2$ $1.3$
66 $245$ $843$ 99 $15$ $0.8$ $1.0$
88 $558$ $2136$ 99 $17$ $2.2$ $3.1$
$10$ $1100$ $4535$ $10$ $16$ $8.4$ $9.5$
$12$ $1963$ $8544$ $11$ $23$ $29.3$ $42.4$
$14$ $3255$ 14 76314\,763 $10$ $26$ $74.6$ $153.3$
$16$ $5100$ 23 88823\,888 99 $29$ $179.9$ $547.3$
$18$ $7638$ 36 71136\,711 $10$ $33$ $504.8$ $1477.5$
2020 11 02511\,025 54 12054\,120 $11$ $28$ $1032.4$ $2735.9$
$22$ 15 43315\,433 77 09977\,099 $10$ $32$ $2318.1$ 13 106.113\,106.1
$24$ 21 05021\,050 106 728106\,728 $11$ $50$ $5811.7$ 36 668.136\,668.1
TABLE IV: Comparison of sequential SOS (S) and coordinate descent (C) for the N-link pendulum ROA example using a bad initial guess.
n Size Iter. [-] Time [s]
$_{\text{con.}} n$$_{\text{dec.}}$ S C S C
44 $85$ $248$ $45$ NaN $3.6$ NaN
66 $245$ $843$ $56$ NaN $7.8$ NaN
88 $558$ $2136$ $73$ NaN $30.9$ NaN
$10$ $1100$ $4535$ $11$ NaN $16.7$ NaN
$12$ $1963$ $8544$ $12$ NaN $52.3$ NaN
$14$ $3255$ 14 76314\,763 $11$ NaN $150.2$ NaN
$16$ $5100$ 23 88823\,888 $12$ NaN $454.2$ NaN
$18$ $7638$ 36 71136\,711 $16$ NaN $1354.01$ NaN
2020 11 02511\,025 54 12054\,120 $15$ NaN $3073.6$ NaN
$22$ 15 43315\,433 77 09977\,099 $14$ NaN $9873.7$ NaN
$24$ 21 05021\,050 106 728106\,728 $15$ NaN 22 245.922\,245.9 NaN
V-C Control Synthesis for Nonlinear Systems
The second case considers the synthesis of a stabilizing control law for nonlinear systems which consider state and control constraints. Besides of directly using such stabilizing control laws, such problems are relevant, for example, to compute terminal conditions needed to ensure asymptotic stability of nonlinear model predictive control. We consider the input constraint set as polyhedral set
$\mathcal{U}=\{u\in\mathbb{R}^{m}~|~H_{\mathcal{U}}u\leq\mathbf{1}_{p}\}$
with $H_{\mathcal{U}}\in\mathbb{R}^{p\times m}$, $p\in\mathbb{N}$, and $\mathbf{1}_{p}=(1,..,1)\in R^{p}$.
We are interested to solve the following SOS problem
$\min_{\begin{subarray}{c}V,\hat{\kappa}\in\mathbb{R}[x],\\ s_{1},s_{2}\in\Sigma[x],\\ s_{3}\in\Sigma[x]^{m}\end{subarray}}$ $\quad\|h-V\|^{2}_{\mathbb{R}[x]}$ (24a)
s.t. $V-\epsilon(x)$ $\in\Sigma[x]$ (24b)
$s_{1}(V-\beta)-\langle\nabla V,f(x,\hat{\kappa})\rangle-\epsilon(x)$ $\in\Sigma[x]$ (24c)
$s_{2}(V-\beta)-h(x)$ $\in\Sigma[x]$ (24d)
$s_{3}(\hat{h}-\beta)-(H_{\mathcal{U}}\hat{\kappa}-\mathbf{1}_{p})$ $\in\Sigma[x]^{m}$ (24e)
where $V(\cdot)$ is a the sought Lyapunov function, $\beta>0$ is the level set (either fixed or an additional decision variable), $\hat{\kappa}$ is the sought polynomial control law, and $s(\cdot)$ are the SOS multiplier resulting from Theorem 1. Equation (24b) and (24c) are standard Lyapunov conditions and (24d) and (24e) enforce state and control constraints, respectively. We make use of a similar cost function as before.
Again two examples are considered: The first example considers non-affine nonlinear dynamics of an aircraft’s longitudinal motion with four states and two controls, which cannot be solved with coordinate-descent in a direct manner. The second showcases the synthesis of a control law for control-affine satellite attitude dynamics.
V-C1 Nonlinear Aircraft Dynamics
In this example, we make use of the four state longitudinal motion of an aircraft provided in [8]. Compared to the previous examples, the control enter non-affine into the dynamics and hence, can not be solved by coordinate descent in a direct manner. For synthesis, we must re-write the system dynamic into an affine form. This would be possible by augmenting the system dynamics with a zero dynamic of the controls, resulting in $n_{e}=n+m$ state and hence in a higher computational load. In contrast, with the sequential approach we are able to directly solve this problem. In Table V, we provide only results for the sequential SOS approach. In about $13\text{\,}\mathrm{s}$ one obtains a safe-by-design controller for the nonlinear system.
V-C2 Satellite Attitude Control Synthesis
[…]
TABLE V: Comparison of sequential SOS (S) and coordinate descent (C) for the four-state constrained ROA problem.
n Size Iter. [-] Time [s] Cost [-]
$_{\text{con.}} n$$_{\text{dec.}}$ S C S C S C
4 1099 7333 20 - 12.5 - 0.34 -
6 5543 57048 5 6 73.9 91.1 0.01 29.6
V-D Reachability Analysis
Our third case deals with the computation of the inner-approximation of the reachable set, using two different dynamical systems, with two and four states respectively. The problem formulation and the derivation can be found in [53, 11]. The primary goal is to synthesize a set in the form $\Omega^{V}_{t,\gamma}=\{x\in\mathbb{R}^{\ell}\mid V(x,t)\leq\gamma\}$, where $V:\mathbb{R}^{\ell}\times\mathbb{R}\rightarrow\mathbb{R}$ is a reachability storage function and $\gamma>0$ bounds the set, together with a time-varying control law $u(t)=\hat{\kappa}(t,x(t))$
We slightly modify the problem by using a similar cost as in the first benchmark, instead of a bisection approach in [53]. We are interested to solve the following SOS problem (cf.[53, 11])
$\min_{\begin{subarray}{c}s_{0}\in\Sigma[x],\beta\in\mathbb{R},\\ V,\hat{\kappa}\in\mathbb{R}[t,x],\\ s_{1},s_{2},s_{3},s_{4}\in\Sigma[x,t],\\ s_{5},s_{6},\in\Sigma[x,t]^{m}\\ \end{subarray}}\quad\|h-V\|^{2}$ (25a)
$\text{s.t.}\quad s_{0}V(T,x)-P$ $\in\Sigma[x]$ (25b)
$s_{1}(V-\beta)-s_{2}g-h$ $\in\Sigma[t,x]$ (25c)
$s_{3}(V-\beta)-s_{4}g-\tau$ $\in\Sigma[t,x]$ (25d)
$s_{5}(V-\beta)-s_{6}g-(H_{\mathcal{U}}\hat{\kappa}-\mathbf{1}_{p})$ $\in\Sigma[t,x]^{m}$ (25e)
[…]
The results are summarized in Table VI. Both algorithms are initialized using the Lyapunov function from the linearized system. For sequential, all multipliers are initialized with simple quadratic polynomials.
TABLE VI: Comparison of sequential SOS (S) and coordinate descent (C) for reachability analysis for different system sizes.
n Size Iter. [-] Time [s] cost [-]
$_{\text{con.}} n$$_{\text{dec.}}$ S C S C S C
2 943 6042 51 78 22.5 21.5 0.84 0.86
4 3171 27647 11 100 42.8 371.2 0.04 0.39
In the first case, sequential needed significantly less iterations. Yet, both approaches needed an equal total time to solve the problem, with CD needing slightly less.
For the 4D example, the sequential SOS outperforms CD. The overall solve time is lower due to the comparable low number of iterations. Sequential was able to solve both problems to a smaller cost. For the second, we were not able to find an optimal sublevel set $\beta$ for coordinate descent. It either turned into infeasibility or was computed to zero by the underlying solver. Thus, for this problem we set $\beta=0.1$ (larger values led to infeasibility), a feasible but potentially conservative, value. The search for the sublevel set in that case could be carried out by an additional bisection which in turn results in a higher computational effort.
V-E Compatible Control Barrier and Control Lyapunov Function Synthesis
The final case addresses the synthesis of compatible Control Barrier Function (CBF) and a Control Lyapunov Function (CLF). Compatible means both conditions are satisfied simultaneously, which is possible by constructing a unified control law that is shared by all CBFs and the CLF [45]. We make use of an adapted version from [45] by also incorporating control constraints and the $\mathcal{K}$ functions. The nonconvex SOS problem reads
[…]
$\text{s.t. }\hat{V}-\varepsilon(x^{\top}x)$ $\in\Sigma[x]$ (26b)
[…]
with $\tau=\langle\nabla\hat{V}(x),f(x,\hat{\kappa}(x))\rangle+\gamma_{V}V$ and where $\gamma_{h}$ and $\gamma_{V}$ are extended $\mathcal{K}$ functions of the form a(s)=asa(s)=a\,s with $a>0$. The problem deals with constrained satellite control, with one CBF and one CLF, using the same parameter as from Subsection V-C2. Unlike the previous cases, we concentrate more on the computational aspects of the proposed sequential SOS approach and do not compare to CD. The problem has 1756 constraints and 9404 decision variables. The proposed sequential SOS approach needed 53 iterations to find an optimal solution with optimal cost $J(\hat{V}^{*},\hat{h}^{*},\hat{\kappa}^{*})=0.04$. A detailed breakdown of the computational effort is provided in Table VII. The largest portion is to solve the underlying quadratic subproblem ($67\%$), followed by the filter line-search ($32\%$). The high computational effort of the line-search can be primarily traced back to the constraint violation check using the signed-distance approach ($\approx 70\%$). Other components such as Hessian approximation or function evaluations are in total $\approx 1\%$ and are negligible compared to the SDPs.
TABLE VII: Timing results for the CBF-CLF Satellite Example in seconds.
Total time (Alg. 3) Prob. (3) Alg. 2 Others
77.7 52 24.7 0.9
V-F Discussion and Practical Consideration
The proposed sequential SOS algorithm with its filter line-search allows to directly deal with nonconvex, nonlinear SOS problems, even with non-control affine polynomial system dynamics. Unlike coordinate descent, there is no requirement that the initial guess is a feasible solution. Coordinate descent has in general no convergence guarantees, but local convergence guarantees can be given for the sequential approach under mild assumptions. If the problem is infeasible, confirmed by the feasibility restoration, the user might use these information to adapt the problem, i.e., by changing the polynomial structure or revise the problem formulation. The feasibility restoration adds robustness to the overall algorithm, as demonstrated, for example, by the N-link robot arm example. However, there are no guarantees for the overall algorithm or feasibility restoration to find a feasible point. Whereas a non-feasible initial guess might be advantageous, a bad initial guess might lead to insufficient progress or even divergence since it is a Newtons-based method. The filter line search alleviates the problem, but does not fully resolve it. Strategies for initial guess generation are needed, some of which are outlined next.
  • If possible, use information from the problem itself, e.g., linearize the system and compute a linear control law and corresponding for Lyapunov function.
  • Use a simple initial guess (e.g., all polynomials are initialized with basis one) and solve only for a few iterations. Take the result and either use it as an initial guess or use the information to construct an initial guess from it.
  • Solve a simplified version, e.g., remove some of the constraints or relax the constraint bounds.
  • Remove cost function and try to solve a feasibility problem and/or relax the thresholds for feasibility.
Whereas sequential SOS is capable of solving nonconvex problems in a more direct manner and can reduce the computational effort, it is still restricted by the capabilities of the underlying SDP solver, i.e., the computational effort can still be high for large problems. Yet, the proposed approach is one step closer to more tractable methods for nonlinear system analysis and control design using SOS.
VI Conclusion
This paper demonstrates significant computational savings for nonconvex sum-of-squares problems arising in system analysis and control design are leveraging a filter line-search algorithm. In several real-world inspired case studies, it is demonstrated that the proposed approach outperforms state-of-the-art methods, thereby making a step towards more practical synthesis and analysis tools for system’s control. Besides, of comprehensive benchmarking, this paper is accommodated with an open-source implementation and provides local convergence analysis.
Acknowledgement
The authors thank Benoît Legat for his input for the paper and Renato Loureiro for his help to improve the manuscript.